the combined ages of a women, her mother, and her sister equal 130. the mother is 1.5 times as much of the age of the women. the sister is 10 years younger than the women. what is the age of each?
Answer by colliefan(242)

the combined ages of a women, her mother, and her sister equal 130. the mother is 1.5 times as much of the age of the women. the sister is 10 years younger than the women. what is the age of each?
The ages can be algebraically represented as:
woman x
sister x-10
mother 1.5(x)
The sum of these is 130 so:
x + x-10 + 1.5x = 130
2x-10+1.5x=130
3.5x-10=130
3.5x-10+10=130+10
3.5x = 140
x= 40


Substitute this value back into the original equation to make sure it is correct.